Suppose that you and a friend are playing cards and you decide to make a friendly wager. The bet is that you will draw two cards without replacement from a standard deck. If both cards are diamonds, your friend will pay you $487. Otherwise, you have to pay your friend $29. Step 1 of 2 : What is the expected value of your bet? Round your answer to two decimal places. Losses must be expressed as negative values.
Answer #1

P(both cards are diamonds)=(13/52)*(12/51)

expected value of your bet =487*(13/52)*(12/51)-29*(1-(13/52)*(12/51))= 1.35

\